Two Step Tariff. Determine the equivalent quota.
D: P = 124 - 2Q
S: P = 44 + 2Q
Tariff: P = 40.
1
Expert's answer
2016-05-26T12:43:03-0400
D: P = 124 - 2Q
S: P = 44 + 2Q.
In equilibrium Qd = Qs and P is the same, so:
124 - 2Q = 44 + 2Q,
4Q = 80,
Q = 20 units.
P = 124 - 2*20 = $84.
If a tariff of P = $40 is imposed, then new price is P = $124, then the new quantity demanded is:
124 = 124 - 2Q,
Qd = 0 units.
The new quantity supplied is:
124 = 44 + 2Q,
2Q = 80,
Qs = 40 units.
So, the equivalent quota is Qs - Qd = 40 - 0 = 40 units.